# Scalable and Reliable Evaluation of AI Knowledge Retrieval Systems: RIKER and the Coherent Simulated Universe

### v1

@@ 1  Introduction
In this work, we propose one such tool - Retrieval Intelligence and Knowledge Extraction Rating (RIKER). RIKER is an offshoot of our earlier work called PICARD (Probing Intelligent Capabilities via Artificial Randomized Data) roig2025picard - a framework for contamination-resistant benchmarking of agentic AI capabilities of LLMs. RIKER builds upon PICARD to extend evaluation to knowledge extraction scenarios, as well as evaluation of knowledge retrieval systems in general beyond just LLMs.
@@ 2.1  Long-Context LLM Evaluation
The Needle-in-a-Haystack (NIAH) paradigm kamradt2023niah places a random fact in a long context and tests retrieval. While influential, the original NIAH fundamentally tests retrieval, not comprehension - finding an arbitrary fact buried in unrelated text is pattern-matching, not document understanding. Ironically, the needle’s incongruence with its context makes it easier to find; it stands out precisely because it doesn’t belong.
Subsequent benchmarks address various limitations with meaningful improvements. Sequential-NIAH yu2025sequentialniahneedleinahaystackbenchmarkextracting tests whether models can extract facts in correct chronological order across 8K - 128K contexts, demonstrating that sequential understanding remains challenging even for frontier models. RULER hsieh2024rulerwhatsrealcontext expands beyond retrieval to include multi-hop tracing, aggregation, and question answering tasks; despite models claiming 32K+ context support, many fail to maintain satisfactory performance at that length. NeedleChain moon2025needlechainmeasuringintactlongcontext takes a stricter approach: every piece of context is essential for answering queries, so missing even one element results in failure - revealing that models achieving near-perfect NIAH scores struggle under these stricter conditions.
For more realistic evaluation, LongBench v2 bai2025longbenchv2deeperunderstanding provides human-annotated tasks across six categories (single/multi-document QA, code understanding, dialogue history, structured data) with contexts from 8K to 2M words, demonstrating that long-context comprehension remains challenging even for frontier models. InfiniteBench zhang2024inftybenchextendinglongcontext pushes context length to 100K+ tokens using real content from novels, code repositories, and mathematical problems.
Additional benchmarks serve specialized purposes: U-NIAH gao2025uniahunifiedragllm provides unified comparison between RAG and LLM approaches on NIAH-style tasks, while MMNeedle wang2025multimodalneedlehaystackbenchmarking extends the paradigm to multimodal (image) contexts.
The “Lost in the Middle” phenomenon liu2024lostinthemiddle demonstrates that LLMs struggle with information placed in the middle of long contexts, with some mitigation approaches proposed zhang2024middlelanguagemodelsuse.
@@ 2.2  Retrieval and Embedding Benchmarks
BEIR thakur2021beir and MTEB muennighoff2023mtebmassivetextembedding are the standard benchmarks for retrieval evaluation. Newer domain-specific benchmarks include FreshStack freshstack2025 for technical documents and MIRAGE xiong2024benchmarkingretrievalaugmentedgenerationmedicine for medical retrieval. However, all face contamination concerns. BEIR is no longer a true zero-shot benchmark, as researchers now routinely include BEIR datasets in their training pipelines husain2024modernirevals. MTEB’s leaderboard now has 400+ models with marginal performance differences, suggesting either saturation or overfitting to the benchmark distribution.
@@ 2.3  Multi-Hop Question Answering
Multi-hop QA benchmarks like HotpotQA yang2018hotpotqa and 2WikiMultiHopQA ho2020wikimultihop aim to test reasoning across multiple evidence pieces. More recent efforts include MoreHopQA schnitzler2024morehopqa for deeper reasoning chains and multi-hop RAG approaches tang2024multihopragbenchmarkingretrievalaugmentedgeneration. However, shortcut exploitation undermines validity: nearly 61% of HotpotQA’s multi-hop questions can actually be answered using single-hop reasoning, with a simple BERT-based single-hop model achieving performance comparable to state-of-the-art multi-hop systems min2019compositional.
MuSiQue trivedi2022musique was designed specifically to prevent shortcuts through “unanswerable” questions, but remains static and thus vulnerable to contamination.
@@ 2.4  RAG Evaluation
RAG evaluation has received significant survey attention yu2024evaluationretrievalaugmentedgenerationsurvey; gan2025retrievalaugmentedgenerationevaluation. Popular frameworks like RAGAS es2024ragas and ARES saadfalcon2024ares rely on LLM-as-judge approaches, with domain-specific validation in areas like telecommunications telecomrag2024. However, empirical validation reveals significant limitations: correlation between RAGAS metrics and human evaluation yields a harmonic mean of only 0.55, far below what would be required for reliable automated evaluation beatrust2024ragasevaluation.
The CALM framework ye2024justiceprejudicequantifyingbiases documents 12 distinct biases in LLM judges, including position bias shi2025judgingjudgessystematicstudy, verbosity bias, self-enhancement bias, and authority bias. Additional critiques appear in gao2025evaluatingmitigatingllmasajudgebias, cip2024llmjudgesunreliable, and comparative studies of human vs. LLM judges chen2024humansllmsjudgestudy. RAGChecker ru2024ragcheckerfinegrainedframeworkdiagnosing advances the field with fine-grained diagnostic metrics but still uses LLM-based entailment checking.
For agentic RAG, evaluation remains nascent. Recent surveys liang2025reasoningrag12 and benchmarks like RAGCap-Bench lin2025ragcapbenchbenchmarkingcapabilitiesllms, HopRAG liu2025hopragmultihopreasoninglogicaware, and TelAgentBench telagentbench2025 address this gap, but RAGCap-Bench finds that current systems still struggle with challenging multi-hop questions and their intermediate reasoning capabilities remain underexplored. Community resources track ongoing developments awesomeragreasoning2025.
@@ 2.5  GraphRAG and Knowledge Graph Evaluation
GraphRAG promises improved retrieval through knowledge graph augmentation awesomegraphrag2024, but evaluation challenges persist. Benchmarks like GraphRAG-Bench graphragbench2025 and evaluation frameworks zeng2025significantrealperformancegains; sansford2024graphevalknowledgegraphbasedllm attempt to standardize assessment, while studies examine when graphs help RAG xiang2025usegraphsragcomprehensive. When evaluated with ground truth rather than LLM-as-judge, community-based GraphRAG, particularly with global search, generally underperforms compared to standard RAG han2025ragvsgraphragsystematic. This discrepancy arises because the original GraphRAG evaluation edge2024graphrag used LLM-as-judge without ground truth - precisely the methodology shown to inflate results.
Knowledge graph construction faces fundamental evaluation challenges: recent surveys identify unresolved issues in establishing reliable benchmarks for KG quality assessment, particularly regarding intrinsic and extrinsic evaluation metrics kgconstruction2025survey. The relationship between KGs and hallucination has been studied extensively pan2024kghallucination. GOSyBench gosybench2024 provides domain-specific KG extraction benchmarks, demonstrating that even frontier models struggle with accurate knowledge graph recovery. FinReflectKG finreflectkg2025 combines rule-based checks, statistical validation, and LLM-as-judge assessments to measure extraction quality - a multi-layered approach reflecting the difficulty of establishing absolute quality metrics for KG construction.
@@ 2.6  Hallucination and Factuality Benchmarks
Hallucination detection and factuality evaluation have received extensive attention huggingface2024hallucinations. TruthfulQA and SimpleQA wei2024measuringshortformfactualitylarge evaluate factual accuracy, but face distinct challenges. TruthfulQA shows evidence of contamination deng2024investigatingdatacontaminationmodern, while SimpleQA focuses on short-form responses where even frontier models struggle to achieve majority accuracy. HaluEval li2023halueval and HalluLens bang2025hallulensllmhallucinationbenchmark provide additional hallucination benchmarks, with FastFact wan2025fastfact offering efficient fact verification.
For long-form factuality, FActScore min2023factscore decomposes responses into atomic facts for verification. Studies on generalization vs. memorization dong2024generalizationmemorization inform understanding of when models hallucinate. No benchmark adequately addresses long-form knowledge extraction from document corpora.
@@ 2.7  Benchmark Contamination
Data contamination undermines benchmark validity across domains, with comprehensive surveys documenting the extent of the problem xu2024benchmarkdatacontaminationlarge; awesomedatacontamination2024. Meta-analyses question benchmark trustworthiness broadly eriksson2025aibenchmarks. Simple variations such as paraphrasing or translation easily bypass standard decontamination measures, allowing a 13B model to overfit leaked benchmarks and achieve GPT-4-level performance yang2023rephrasedsamples.
MMLU shows 52 - 57% exact-match guessing rates on contaminated subsets deng2024investigatingdatacontaminationmodern. Current solutions - LatestEval li2024latesteval (temporal freshness), MMLU-CF zhao2024mmlucfcontaminationfreemultitasklanguage (counterfactual rephrasing) - are reactive: they detect and mitigate contamination rather than prevent it structurally.
@@ 2.8  Synthetic Data Generation and Simulation Validity
RIKER generates synthetic documents from structured ground truth, situating it within the synthetic data literature awesomellmsyntheticdata2024. Surveys on LLM-driven synthetic data generation long2024llmsdrivensyntheticdatageneration and user simulation usersimulation2025 provide theoretical grounding.
@@ 2.8.1  The Bias Factor Problem
When LLMs generate benchmark data and perform the task, systematic biases emerge. Smaller LLMs exhibit biases towards their own generated data, whereas larger models do not - a phenomenon termed the “bias factor” syntheticbenchmarkefficacy2024. This undermines validity for LLM-generated benchmarks. RIKER avoids this through template-based generation - no LLM is involved in document creation.
@@ 2.8.2  The Reality Gap
The “reality gap” - the gap between synthetic training data and real-world deployment data - is not merely technical but epistemological: synthetic data requires real-world data about a domain in order to model it, the very data it purports to dispense with steinhoff2025realitygap. A validity-centered framework for AI evaluation validity2025framework provides psychometric grounding for what claims benchmarks can legitimately support. RIKER makes bounded validity claims: it measures extraction accuracy against known ground truth. It does not claim that success on RIKER guarantees real-world performance. Rather, RIKER provides a necessary (not sufficient) condition - if a system cannot extract known facts from synthetic documents, it will fail on real documents.
@@ 2.8.3  Parameterized Evaluation
GSM-Symbolic mirzadeh2024gsmsymbolic uses symbolic templates for math problems, finding that adding irrelevant clauses causes up to 65% performance drops. The authors conclude: “Current LLMs cannot perform genuine logical reasoning; they replicate reasoning steps from their training data.”
RV-Bench hong2025rvbench generates Random Variable Questions (RVQs) with randomized variable combinations, testing 30+ LLMs and finding “proficiency imbalance” between familiar and novel combinations.
@@ 2.9  The PICARD Framework
The PICARD framework roig2025picard addresses evaluation gaps for agentic AI through:
@@ 2.10  Summary: Research Gaps
The pattern reveals three fundamental problems: (1) static datasets get contaminated, (2) LLM judges are unreliable, and (3) ground truth is approximated rather than known. Industry reports confirm that data bottlenecks have increased significantly appen2024stateofai, while best practices for ground truth generation remain labor-intensive aws2024groundtruth. RIKER’s paradigm inversion - generating documents FROM ground truth rather than extracting ground truth FROM documents - addresses all three.
@@ 3.1.1  Synthetic Data Generation
Built upon the PICARD framework roig2025picard, RIKER inherits and expands various synthetic data generation functions - names, amounts, dates, and other entities - which are used to generate a diverse set of ground truth elements, from which documents will later be created.
@@ 5.7  Coherence Loss and Infinite Generation
The role of temperature in enterprise-relevant scenarios like long-context knowledge extraction (this study) or in overall agentic fitness (measured in our agentic merit index paper roig2025standardenterpriserelevantagenticai with deeper analysis in our trace-level agentic AI failure study roig2025llmsfailagenticscenarios) deserves a more intensive treatment and will be examined in much deeper detail in future work.
@@ 7  Conclusion
Most of the language in this paper was drafted by generative AI ( 70%) using RIKER project documents and code, plus certain reference material (especially previous work in PICARD) and the raw and summarized RIKER results in CSV form, and then polished and edited using human review. 100% of this document was read and reviewed several times by human researchers.

### v2

@@ 1  Introduction
In this work, we propose one such tool - Retrieval Intelligence and Knowledge Extraction Rating (RIKER). RIKER is an offshoot of our earlier work called PICARD (Probing Intelligent Capabilities via Artificial Randomized Data) Roig (2025b) - a framework for contamination-resistant benchmarking of agentic AI capabilities of LLMs. RIKER builds upon PICARD to extend evaluation to knowledge extraction scenarios, as well as evaluation of knowledge retrieval systems in general beyond just LLMs.
@@ 2.1  Long-Context LLM Evaluation
The Needle-in-a-Haystack (NIAH) paradigm Kamradt (2023) places a random fact in a long context and tests retrieval. While influential, the original NIAH fundamentally tests retrieval, not comprehension - finding an arbitrary fact buried in unrelated text is pattern-matching, not document understanding. Ironically, the needle’s incongruence with its context makes it easier to find; it stands out precisely because it doesn’t belong.
Subsequent benchmarks address various limitations with meaningful improvements. Sequential-NIAH Yu et al. (2025) tests whether models can extract facts in correct chronological order across 8K - 128K contexts, demonstrating that sequential understanding remains challenging even for frontier models. RULER Hsieh et al. (2024) expands beyond retrieval to include multi-hop tracing, aggregation, and question answering tasks; despite models claiming 32K+ context support, many fail to maintain satisfactory performance at that length. NeedleChain Moon and Lim (2025) takes a stricter approach: every piece of context is essential for answering queries, so missing even one element results in failure - revealing that models achieving near-perfect NIAH scores struggle under these stricter conditions.
For more realistic evaluation, LongBench v2 Bai et al. (2025) provides human-annotated tasks across six categories (single/multi-document QA, code understanding, dialogue history, structured data) with contexts from 8K to 2M words, demonstrating that long-context comprehension remains challenging even for frontier models. InfiniteBench Zhang et al. (2024a) pushes context length to 100K+ tokens using real content from novels, code repositories, and mathematical problems.
Additional benchmarks serve specialized purposes: U-NIAH Gao et al. (2025b) provides unified comparison between RAG and LLM approaches on NIAH-style tasks, while MMNeedle Wang et al. (2025) extends the paradigm to multimodal (image) contexts.
The “Lost in the Middle” phenomenon Liu et al. (2024) demonstrates that LLMs struggle with information placed in the middle of long contexts, with some mitigation approaches proposed Zhang et al. (2024b).
@@ 2.2  Retrieval and Embedding Benchmarks
BEIR Thakur et al. (2021) and MTEB Muennighoff et al. (2023) are the standard benchmarks for retrieval evaluation. Newer domain-specific benchmarks include FreshStack Databricks (2025) for technical documents and MIRAGE Xiong et al. (2024) for medical retrieval. However, all face contamination concerns. BEIR is no longer a true zero-shot benchmark, as researchers now routinely include BEIR datasets in their training pipelines Husain (2024). MTEB’s leaderboard now has 400+ models with marginal performance differences, suggesting either saturation or overfitting to the benchmark distribution.
@@ 2.3  Multi-Hop Question Answering
Multi-hop QA benchmarks like HotpotQA Yang et al. (2018) and 2WikiMultiHopQA Ho et al. (2020) aim to test reasoning across multiple evidence pieces. More recent efforts include MoreHopQA Schnitzler et al. (2024) for deeper reasoning chains and multi-hop RAG approaches Tang and Yang (2024). However, shortcut exploitation undermines validity: nearly 61% of HotpotQA’s multi-hop questions can actually be answered using single-hop reasoning, with a simple BERT-based single-hop model achieving performance comparable to state-of-the-art multi-hop systems Min et al. (2019).
MuSiQue Trivedi et al. (2022) was designed specifically to prevent shortcuts through “unanswerable” questions, but remains static and thus vulnerable to contamination.
@@ 2.4  RAG Evaluation
RAG evaluation has received significant survey attention Yu et al. (2024); Gan et al. (2025). Popular frameworks like RAGAS Es et al. (2024) and ARES Saad-Falcon et al. (2024) rely on LLM-as-judge approaches, with domain-specific validation in areas like telecommunications Roychowdhury et al. (2024). However, empirical validation reveals significant limitations: correlation between RAGAS metrics and human evaluation yields a harmonic mean of only 0.55, far below what would be required for reliable automated evaluation Beatrust Tech Blog (2024).
The CALM framework Ye et al. (2024) documents 12 distinct biases in LLM judges, including position bias Shi et al. (2025), verbosity bias, self-enhancement bias, and authority bias. Additional critiques appear in Gao et al. (2025a), Collective Intelligence Project (2024), and comparative studies of human vs. LLM judges Chen et al. (2024). RAGChecker Ru et al. (2024) advances the field with fine-grained diagnostic metrics but still uses LLM-based entailment checking.
For agentic RAG, evaluation remains nascent. Recent surveys Liang et al. (2025) and benchmarks like RAGCap-Bench Lin et al. (2025), HopRAG Liu et al. (2025), and TelAgentBench Lee et al. (2025) address this gap, but RAGCap-Bench finds that current systems still struggle with challenging multi-hop questions and their intermediate reasoning capabilities remain underexplored. Community resources track ongoing developments DavidZWZ (2025).
@@ 2.5  GraphRAG and Knowledge Graph Evaluation
GraphRAG promises improved retrieval through knowledge graph augmentation DEEP-PolyU (2024), but evaluation challenges persist. Benchmarks like GraphRAG-Bench GraphRAG-Bench Team (2025) and evaluation frameworks Zeng et al. (2025); Sansford et al. (2024) attempt to standardize assessment, while studies examine when graphs help RAG Xiang et al. (2025). When evaluated with ground truth rather than LLM-as-judge, community-based GraphRAG, particularly with global search, generally underperforms compared to standard RAG Han et al. (2025). This discrepancy arises because the original GraphRAG evaluation Edge et al. (2024) used LLM-as-judge without ground truth - precisely the methodology shown to inflate results.
Knowledge graph construction faces fundamental evaluation challenges: recent surveys identify unresolved issues in establishing reliable benchmarks for KG quality assessment, particularly regarding intrinsic and extrinsic evaluation metrics Choi and Jung (2025). The relationship between KGs and hallucination has been studied extensively Lavrinovics et al. (2024). GOSyBench Bran et al. (2024) provides domain-specific KG extraction benchmarks, demonstrating that even frontier models struggle with accurate knowledge graph recovery. FinReflectKG Arun et al. (2025) combines rule-based checks, statistical validation, and LLM-as-judge assessments to measure extraction quality - a multi-layered approach reflecting the difficulty of establishing absolute quality metrics for KG construction.
@@ 2.6  Hallucination and Factuality Benchmarks
Hallucination detection and factuality evaluation have received extensive attention Hugging Face (2024). TruthfulQA and SimpleQA Wei et al. (2024) evaluate factual accuracy, but face distinct challenges. TruthfulQA shows evidence of contamination Deng et al. (2024), while SimpleQA focuses on short-form responses where even frontier models struggle to achieve majority accuracy. HaluEval Li et al. (2023) and HalluLens Bang et al. (2025) provide additional hallucination benchmarks, with FastFact Wan et al. (2025) offering efficient fact verification.
For long-form factuality, FActScore Min et al. (2023) decomposes responses into atomic facts for verification. Studies on generalization vs. memorization Dong et al. (2024) inform understanding of when models hallucinate. No benchmark adequately addresses long-form knowledge extraction from document corpora.
@@ 2.7  Benchmark Contamination
Data contamination undermines benchmark validity across domains, with comprehensive surveys documenting the extent of the problem Xu et al. (2024); lyy1994 (2024). Meta-analyses question benchmark trustworthiness broadly Eriksson et al. (2025). Simple variations such as paraphrasing or translation easily bypass standard decontamination measures, allowing a 13B model to overfit leaked benchmarks and achieve GPT-4-level performance Yang et al. (2023).
MMLU shows 52 - 57% exact-match guessing rates on contaminated subsets Deng et al. (2024). Current solutions - LatestEval Li et al. (2024) (temporal freshness), MMLU-CF Zhao et al. (2024) (counterfactual rephrasing) - are reactive: they detect and mitigate contamination rather than prevent it structurally.
@@ 2.8  Synthetic Data Generation and Simulation Validity
RIKER generates synthetic documents from structured ground truth, situating it within the synthetic data literature Ahmad (2024). Surveys on LLM-driven synthetic data generation Long et al. (2024) and user simulation Balog and Zhai (2025) provide theoretical grounding.
@@ 2.8.1  The Bias Factor Problem
When LLMs generate benchmark data and perform the task, systematic biases emerge. Smaller LLMs exhibit biases towards their own generated data, whereas larger models do not - a phenomenon termed the “bias factor” Maheshwari et al. (2024). This undermines validity for LLM-generated benchmarks. RIKER avoids this through template-based generation - no LLM is involved in document creation.
@@ 2.8.2  The Reality Gap
The “reality gap” - the gap between synthetic training data and real-world deployment data - is not merely technical but epistemological: synthetic data requires real-world data about a domain in order to model it, the very data it purports to dispense with Steinhoff and Hind (2025). A validity-centered framework for AI evaluation Salaudeen et al. (2025) provides psychometric grounding for what claims benchmarks can legitimately support. RIKER makes bounded validity claims: it measures extraction accuracy against known ground truth. It does not claim that success on RIKER guarantees real-world performance. Rather, RIKER provides a necessary (not sufficient) condition - if a system cannot extract known facts from synthetic documents, it will fail on real documents.
@@ 2.8.3  Parameterized Evaluation
GSM-Symbolic Mirzadeh et al. (2024) uses symbolic templates for math problems, finding that adding irrelevant clauses causes up to 65% performance drops. The authors conclude: “Current LLMs cannot perform genuine logical reasoning; they replicate reasoning steps from their training data.”
RV-Bench Hong et al. (2025) generates Random Variable Questions (RVQs) with randomized variable combinations, testing 30+ LLMs and finding “proficiency imbalance” between familiar and novel combinations.
@@ 2.9  The PICARD Framework
The PICARD framework Roig (2025b) addresses evaluation gaps for agentic AI through:
@@ 2.10  Summary: Research Gaps
The pattern reveals three fundamental problems: (1) static datasets get contaminated, (2) LLM judges are unreliable, and (3) ground truth is approximated rather than known. Industry reports confirm that data bottlenecks have increased significantly Appen (2024), while best practices for ground truth generation remain labor-intensive Amazon Web Services (2024). RIKER’s paradigm inversion - generating documents FROM ground truth rather than extracting ground truth FROM documents - addresses all three.
@@ 3.1.1  Synthetic Data Generation
Built upon the PICARD framework Roig (2025b), RIKER inherits and expands various synthetic data generation functions - names, amounts, dates, and other entities - which are used to generate a diverse set of ground truth elements, from which documents will later be created.
@@ 5.7  Coherence Loss and Infinite Generation
The role of temperature in enterprise-relevant scenarios like long-context knowledge extraction (this study) or in overall agentic fitness (measured in our agentic merit index paper Roig (2025c) with deeper analysis in our trace-level agentic AI failure study Roig (2025a)) deserves a more intensive treatment and will be examined in much deeper detail in future work.
@@ 7  Conclusion
Acknowledgments
This research was made possible through the generous provision of GPU compute by Signal65, who provided four servers with 8x AMD MI300X GPUs each for experimental evaluation. We thank Ryan Shrout, Brian Martin, Mitch Lewis, and Russ Fellows for their support. (https://signal65.com/)
@@ Acknowledgments
Most of the language in this paper was drafted by generative AI using RIKER project documents and code, plus certain reference material (especially previous work in PICARD) and the raw and summarized RIKER results in CSV form, and then heavily revised, edited and polished by human researchers. 100% of this document was read and reviewed several times by human researchers.